OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -4 2x - y = 1

OpenStudy (anonymous):

x=-4+2y 2(-4+2y)-y=1 -8+4y-y=1 -8+3y=1 3y=1--8 3y=9 y=3 x-2y=-4 x-2(3)=-4 x-6=-4 x=-4+6 x=2 (2,3)